# Multi-bandpass Photometry for Exoplanet Atmosphere Reconnaissance (MPEAR) with the Habitable Worlds Observatory (HWO) I. Differentiating Earth from Neptunes During Discovery

### v1

@@ 5 Noise Modeling
Having identified the cold and warm Neptunes that could reasonably be in an orbital configuration consistent with the detected planet (condition 1, Section 3), and whose flux at 500 nm is similar to that of the Earth at quadrature (condition 2, Section 4), we now investigate whether additional photometric information can distinguish between these scenarios. This requires modeling the expected noise over the full wavelength range.
@@
pyEDITH  was originally adapted from the ETC within the Altruistic Yield Optimization code (AYO, 2014ApJ...795..122S), which is currently being used for yield estimates in the HWO community. The ETC module of AYO, written in IDL and C, has been released open-source [^6] https://starkspace.com. As part of our pyEDITH  efforts, we first translated the code into Python, then we implemented the seamless ingestion of both the EAC YAML files [^7] https://github.com/HWO-GOMAP-Working-Groups/Sci-Eng-Interface/tree/main/hwo_sci_eng and the coronagraph Yield Input Packages (YIP) files that are produced by different teams developing coronagraph designs and technologies for HWO. We describe in detail how to use pyEDITH and the noise terms that are considered for the exposure time calculation in pyedith2025 and in the online documentation [^8] https://pyedith.readthedocs.io/en/latest/index.html.
@@
In Table 2 we list the parameters used in the noise simulation; these are the same for all simulated scenarios. In our simulations, we used a simplified observatory through the special “Toy Model” class in pyEDITH. The “Toy Model” observatory is a simplified version of the first Exploratory Analysis Case (EAC1) considered for HWO. The mirror dimensions follow the prescribed ones for the EAC1 case, but all throughputs and coronagraph specifications are averaged over wavelength, so that every value at every wavelength is constant. This is an approximation, but it is a valuable one, since it allows us to provide the community with a more generalized answer that would inform future studies on coronagraph and detector specifications, which we will also explore in a future study (see Section 7.4). A Toy Model observatory was also used in the validation of other existing ETCs for HWO in 2025arXiv250218556S.
We consider the Earth at quadrature and the best warm and cold Neptunes identified in Section 4. We first calculate the exposure time $\tau_{discovery}$ that would correspond to S/NS/N=7 at the discovery bandpass (centered at $\lambda_{discovery}=0.5\ \mu m$ and with a 20% bandwidth) for the three planets in our study. This value is $\approx 3.17\ hr$. As a reminder, all planets have comparable $\tau_{discovery}$ by construction, as we purposely chose the Neptune planets whose flux in the $\lambda_{discovery}$ bandpass was the closest to the Earth’s flux (see Section 4).
@@
This approach has been widely used in literature to determine the detectability of a spectral feature, or as a way to discriminate between two models (see e.g., LustigYaeger2019; Bixel2021; 2024arXiv240108492A). We would consider the two models qualitatively distinguishable in a specific bandpass if $S\geq 3$ for that bandpass, and reliably distinguishable if $S\geq 5$. This means that the difference between the two models is greater than three or five times the uncertainty.
5.1 Results
@@
This is especially visible in the cold Neptune scenario, which nearly doubles in significance at its peak around 850 nm when quadrupling the bandwidth (5% to 20%, from $S\sim 1.8$ to $\sim 3.2$ at the peak). At wavelengths between 790 and 850 nm at 20% bandwidth, the cold Neptune can be more easily differentiated from the Earth compared to the warm Neptune case, with $S>3$. The strong significance at this wavelength range reflects the methane and cloud absorption features that distinguish Neptune from the Earth. Observing with one or more secondary bandpass(es) in this range would effectively differentiate the two models (see Section 6).
The warm Neptune is overall harder to differentiate from the Earth, with $S\sim 1.5$ at best independent of the bandwidth. However, for the 20% bandwidth case more wavelength regions across the whole range achieve higher $S$ values, as smaller peaks in significance join to form a larger one. The limited discrimination potential stems from similarities in the two spectra (see Section 4). Longer observing times would increase the $S/N$ and might allow for an easier differentiation, which otherwise is more challenging compared to the cold Neptune case (see Section 7.2).
@@
6 Optimizing Parallel Channel Positions
@@
We identify the best parallel bandpass position for each scenario (Earth versus Warm Neptune, Earth versus Cold Neptune) as the ones that maximize $S$ for each channel (UV, VIS, NIR) and for every simulated bandwidth (5%, 10%, 20%). To do this, we require that the photometric bin is fully included within a single channel and does not overlap with the discovery bandpass. This choice is motivated by technology considerations: 1) two bandpasses cannot overlap during the same observation, and 2) different channels would feed into different coronagraphs and detectors, so there should be a hard cutoff that separates each channel from the others. We assume idealized filters with perfect transmission between the edges of the bandpass and no transmission outside of that range, and that we can freely select the central wavelength. This will surely not be the case in reality, where specific detector responses must be accounted for, and a discrete set of filters will eventually be selected. We discuss the implications of these assumptions in Section 7.4.
@@
In the case where three simultaneous bandpasses are explored, we can also compare the difference in the photometric flux observed between two different pairs of bandpasses. For every flavor of Option 3, we can select the three bandpasses (1, 2, 31,\ 2,\ 3) and calculate the differences in the fluxes that would be acquired in each bandpass, each with its own uncertainty. We can then compare $(F_{1}/F_{2})$ versus $(F_{3}/F_{2})$, as these values represent the difference in color between pairs of bandpass. This is done in a similar fashion in previous literature studies (see e.g. JKT2016; 2018AJ....156..158B; Smith2020). The three bandpasses are not necessarily ordered by increasing wavelength, but we compare each secondary bandpass with respect to the discovery bandpass: $F_{2}$ is centered at 500 nm, while the $F_{1}$ and $F_{3}$ are centered on the secondary bandpass, regardless of where they are located.
@@ 6.1 Results
By comparing the selected bandpasses with the overall behavior of the significance (Figure 6), we notice that there are specific local maxima which often coincide with the best bandpass selected for that channel. For the warm and the cold Neptune, UV peaks at 350 and 300 nm respectively for all bandpasses; in the VIS range, wavelengths around 800-880 nm are preferred; in the near-infrared, the significance $S$ peaks around 1-1.2 $\mathrm{\mu m}$. There is a noticeable exception for the warm Neptune scenario: a strong peak in $S$ around 1.03 micron cannot be leveraged by the 10%- and 20%-wide bandpasses, as the width of the bandpass would intersect the hard threshold wavelength that is defined to separate the VIS and NIR channels. A different cutoff between VIS and NIR could provide enhanced differentiability between the Earth and warm Neptune (see Section 7.3). Critically, the inability to leverage the higher $S$ values at longer VIS wavelengths constrains the VIS channel bandpass optimization for the Warm Neptune scenario to $\sim 640$ nm, which has lower significance and also is not in agreement with the corresponding VIS bandpass for the cold Neptune case. Around 640 nm (i.e., the preferred wavelength for the warm Neptune case at 20% bandwidth) the cold Neptune scenario shows the lowest significance of all the VIS channel bandpass positions, so it is hard to differentiate the cold Neptune with respect to the Earth; on the other hand, it would be much easier to differentiate cold Neptune and Earth at 850-880 nm (highest $S$) but we are faced with a local minimum in $S$ for the warm Neptune scenario. This shows that it might be difficult to find a single bandpass combination (or observation strategy) that is convenient to differentiate Earth-like atmospheres from both warm and cold Neptunes.
In Figure 8 we show the significance of the various configurations of the two and three simultaneous bandpass scenarios (Options 2 and 3), for varying bandwidths. We notice a general increase in $S_{option}$ by considering two or three simultaneous bandpasses compared to the discovery bandpass only (D). This is by construction, since we selected the warm and cold Neptune models based on the smallest difference in flux in the discovery bandpass. The significance is very low in this case, but not zero, because of the uncertainty on the simulated measurement (condition 2, see Section 4).
@@
In the Warm Neptune case none of the options produce $S>3$, meaning that the difference in flux between the Warm Neptune and the Earth is less than three times the corresponding noise (as described in 5). In the two-bandpass scenarios, Option 2a is generally preferred to the others for $\geq 10\%$ bandwidth ($S_{2a,10\%}=1.72$$S_{2a,20\%}=2.20$). When three bandpasses are available, Options 3a (UV+D+NIR) and 3c (UV+D+VIS) at 20% bandwidth are almost equally preferred, with $S\sim 2.7$.
@@
Disregarding for now the uncertainty associated with this measurement, we find that the planets are more effectively separated using narrow bandpasses: for the same option, the value of $d$ increases for decreasing bandwidths, resulting in points wider apart in color-color space. This has been observed also in JKT2016 for a similar study (see Section 7.3 for a more in depth comparison). However, the uncertainty on the flux measurements also becomes increasingly larger at smaller bandwidths (see Figure 6), which then factors into $\Delta d$. This often results in values of $r$ that are overall lower for smaller bandwidths. Exceptions arise when the bandpass combination changes with changes in bandwidth. We discuss this further in Section 7.2.
@@
From this analysis, we find that the combination of 360 nm + 500 nm + 1110 nm would be the most promising to differentiate both the cold and the warm Neptune case. However, the exposure time would need to be increased to reduce the uncertainties and obtain a value of $r\geq 3$; we discuss this in detail in Section 7.2.
7 Discussion
In this work, we assessed the potential of simultaneous multi-bandpass photometry as a key to breaking degeneracies inherent in single-bandpass, single-epoch observations of exoplanets. As shown in Sections 3 and 4, a single detection point can be explained by multiple orbital configurations, and planets with different compositions can produce similar fluxes in the single bandpass that HWO could use for its discovery phase (500 nm, 20% bandwidth). While longer integration times and improvements in the instrumentation can reduce uncertainties, they cannot fully eliminate this fundamental ambiguity.
Visiting the system at multiple epochs is the primary way to assess orbital parameters, and is therefore considered as a core part of HWO’s exoplanet survey strategy and taken into account when calculating exoplanet yields for the mission (see e.g., 2014ApJ...795..122S; 2024SPIE13092E..5MM). Such a strategy, however, comes with relevant complications: as time passes between each visit, the orbital position changes and therefore the portion of illuminated surface, the orientation of the planet, and the atmospheric dynamics would most likely have impacted the spectrum of the planet. Hence, multi-epoch observations are not trivial to interpret, and future work is still needed to model such variations, as well as accurately constrain the atmospheric composition. This is even more complicated for multi-planetary systems, where the pairing of the photometric detection over multiple epochs with the correct planet can be challenging and requires specialized algorithms (see e.g., 2021SPIE11823E..0FM; 2025AAS...24611604H).
@@ 7 Discussion
7.1 Optimizing Bandpass Choices for Planet Differentiation
By using the Exposure Time Calculator that we developed (Section 5, pyedith2025) we estimate the measurement uncertainty for any potential ancillary bandpass given the observing time required to achieve S/NS/N=7 in the discovery observation (500 nm), the current baseline assumed by HWO. This allows us to identify the most promising bandpasses for various combinations of parallel channels to differentiate the Earth from “false positives” such as warm and cold Neptunes, thus informing an observational strategy for HWO optimized for multiple bandpasses (Section 6).
@@ 7.1 Optimizing Bandpass Choices for Planet Differentiation
This fundamental trade-off prevents the bandpass combinations from being universally optimal and highlights the model-dependent nature of our findings - we should expect that different models would yield different results and that a photometric observation strategy might take into account some additional population-related statistics (see Section 7.4).
@@
7.2 $S/N$ Considerations
Figure 11: Total significance at varying $S/N$ values for the 2- (top row) and 3-bandpass options (bottom row) that were identified as promising to differentiate a warm or a cold Neptune from the Earth (see Section 6). In orange, solid lines the significance for the warm Neptune case; in aqua, dashed lines the significance for the cold Neptune case. As dashed lines, the $S_{option}=3$ and $S_{option}=5$ significance thresholds. Labeled in each subplots are the corresponding discovery exposure times (in hours) that are necessary to achieve $S_{option}=3$ and $S_{option}=5$, as calculated with pyEDITH. At the top of each subset, the option keyword, the corresponding bandwidth (BW) and the center wavelengths of the bandpasses considered in the option.
For the bandpass combinations that were identified as promising for discriminating the Earth from either Neptune-like atmosphere at the baseline discovery S/NS/N=7, we ran ancillary noise simulations assuming increasing discovery $S/N$ values. This is necessary to identify the bandpass combinations that would be best suited to qualitatively differentiate multiple planet types whenever possible, since we will not know a priori the nature of the planet when observing it for the first time. Hence, we analyzed how bandpasses that are optimized for a specific case would behave for the opposite scenario, for increasing observing time. This ancillary study provides information on the value that observing for longer exposure times has on the overall result over multiple scenarios, and allows us to prioritize specific bandpass combinations that favor both models at the same time, rather than a single scenario. We discuss future plans on how to generalize this effort in Section 7.4.
In Figure 11 we show the option significance that the two- and three-bandpass options we mentioned in the previous section yield for increasing discovery $S/N$ to differentiate both scenarios. We also label the exposure time that would be necessary to reach $S_{option}=3$ and $S_{option}=5$ significance levels for each model, calculated with pyEDITH. These times also include overheads to slew the telescope and dig the dark hole (see Section 5).

### v2

@@ V Noise Modeling
Having identified the cold and warm Neptunes that could reasonably be in an orbital configuration consistent with the detected planet (condition 1, Section III), and whose flux at 500 nm is similar to that of the Earth at quadrature (condition 2, Section IV), we now investigate whether additional photometric information can distinguish between these scenarios. This requires modeling the expected noise over the full wavelength range.
@@
pyEDITH  was originally adapted from the ETC within the Altruistic Yield Optimization code (AYO, Stark et al., 2014), which is currently being used for yield estimates in the HWO community. The ETC module of AYO, written in IDL and C, has been released open-source [^6] https://starkspace.com. As part of our pyEDITH  efforts, we first translated the code into Python, then we implemented the seamless ingestion of both the EAC YAML files [^7] https://github.com/HWO-GOMAP-Working-Groups/Sci-Eng-Interface/tree/main/hwo_sci_eng and the coronagraph Yield Input Packages (YIP) files that are produced by different teams developing coronagraph designs and technologies for HWO. We describe in detail how to use pyEDITH and the noise terms that are considered for the exposure time calculation in Alei et al. (2025b, a) and in the online documentation [^8] https://pyedith.readthedocs.io/en/latest/index.html.
@@
In Table 2 we list the parameters used in the noise simulation; these are the same for all simulated scenarios. In our simulations, we used a simplified observatory through the special “Toy Model” class in pyEDITH. The “Toy Model” observatory is a simplified version of the first Exploratory Analysis Case (EAC1) considered for HWO. The mirror dimensions follow the prescribed ones for the EAC1 case, but all throughputs and coronagraph specifications are averaged over wavelength, so that every value at every wavelength is constant. This is an approximation, but it is a valuable one, since it allows us to provide the community with a more generalized answer that would inform future studies on coronagraph and detector specifications, which we will also explore in a future study (see Section VII.4). A Toy Model observatory was also used in the validation of other existing ETCs for HWO in Stark et al. (2025).
We consider the Earth at quadrature and the best warm and cold Neptunes identified in Section IV. We first calculate the exposure time $\tau_{discovery}$ that would correspond to S/NS/N=7 at the discovery bandpass (centered at $\lambda_{discovery}=0.5\ \mu m$ and with a 20% bandwidth) for the three planets in our study. This value is $\approx 3.17\ hr$. As a reminder, all planets have comparable $\tau_{discovery}$ by construction, as we purposely chose the Neptune planets whose flux in the $\lambda_{discovery}$ bandpass was the closest to the Earth’s flux (see Section IV).
@@
This approach has been widely used in literature to determine the detectability of a spectral feature, or as a way to discriminate between two models (see e.g., Lustig-Yaeger et al., 2019; Bixel and Apai, 2021; Angerhausen et al., 2024). We would consider the two models qualitatively distinguishable in a specific bandpass if $S\geq 3$ for that bandpass, and reliably distinguishable if $S\geq 5$. This means that the difference between the two models is greater than three or five times the uncertainty.
V.1 Results
@@
This is especially visible in the cold Neptune scenario, which nearly doubles in significance at its peak around 850 nm when quadrupling the bandwidth (5% to 20%, from $S\sim 1.8$ to $\sim 3.2$ at the peak). At wavelengths between 790 and 850 nm at 20% bandwidth, the cold Neptune can be more easily differentiated from the Earth compared to the warm Neptune case, with $S>3$. The strong significance at this wavelength range reflects the methane and cloud absorption features that distinguish Neptune from the Earth. Observing with one or more secondary bandpass(es) in this range would effectively differentiate the two models (see Section VI).
The warm Neptune is overall harder to differentiate from the Earth, with $S\sim 1.5$ at best independent of the bandwidth. However, for the 20% bandwidth case more wavelength regions across the whole range achieve higher $S$ values, as smaller peaks in significance join to form a larger one. The limited discrimination potential stems from similarities in the two spectra (see Section IV). Longer observing times would increase the $S/N$ and might allow for an easier differentiation, which otherwise is more challenging compared to the cold Neptune case (see Section VII.2).
@@
VI Optimizing Parallel Channel Positions
@@
We identify the best parallel bandpass position for each scenario (Earth versus Warm Neptune, Earth versus Cold Neptune) as the ones that maximize $S$ for each channel (UV, VIS, NIR) and for every simulated bandwidth (5%, 10%, 20%). To do this, we require that the photometric bin is fully included within a single channel and does not overlap with the discovery bandpass. This choice is motivated by technology considerations: 1) two bandpasses cannot overlap during the same observation, and 2) different channels would feed into different coronagraphs and detectors, so there should be a hard cutoff that separates each channel from the others. We assume idealized filters with perfect transmission between the edges of the bandpass and no transmission outside of that range, and that we can freely select the central wavelength. This will surely not be the case in reality, where specific detector responses must be accounted for, and a discrete set of filters will eventually be selected. We discuss the implications of these assumptions in Section VII.4.
@@
In the case where three simultaneous bandpasses are explored, we can also compare the difference in the photometric flux observed between two different pairs of bandpasses. For every flavor of Option 3, we can select the three bandpasses (1, 2, 31,\ 2,\ 3) and calculate the differences in the fluxes that would be acquired in each bandpass, each with its own uncertainty. We can then compare $(F_{1}/F_{2})$ versus $(F_{3}/F_{2})$, as these values represent the difference in color between pairs of bandpass. This is done in a similar fashion in previous literature studies (see e.g. Krissansen-Totton et al., 2016; Batalha et al., 2018; Smith et al., 2020). The three bandpasses are not necessarily ordered by increasing wavelength, but we compare each secondary bandpass with respect to the discovery bandpass: $F_{2}$ is centered at 500 nm, while the $F_{1}$ and $F_{3}$ are centered on the secondary bandpass, regardless of where they are located.
@@ VI.1 Results
By comparing the selected bandpasses with the overall behavior of the significance (Figure 6), we notice that there are specific local maxima which often coincide with the best bandpass selected for that channel. For the warm and the cold Neptune, UV peaks at 350 and 300 nm respectively for all bandpasses; in the VIS range, wavelengths around 800-880 nm are preferred; in the near-infrared, the significance $S$ peaks around 1-1.2 $\mathrm{\mu m}$. There is a noticeable exception for the warm Neptune scenario: a strong peak in $S$ around 1.03 micron cannot be leveraged by the 10%- and 20%-wide bandpasses, as the width of the bandpass would intersect the hard threshold wavelength that is defined to separate the VIS and NIR channels. A different cutoff between VIS and NIR could provide enhanced differentiability between the Earth and warm Neptune (see Section VII.3). Critically, the inability to leverage the higher $S$ values at longer VIS wavelengths constrains the VIS channel bandpass optimization for the Warm Neptune scenario to $\sim 640$ nm, which has lower significance and also is not in agreement with the corresponding VIS bandpass for the cold Neptune case. Around 640 nm (i.e., the preferred wavelength for the warm Neptune case at 20% bandwidth) the cold Neptune scenario shows the lowest significance of all the VIS channel bandpass positions, so it is hard to differentiate the cold Neptune with respect to the Earth; on the other hand, it would be much easier to differentiate cold Neptune and Earth at 850-880 nm (highest $S$) but we are faced with a local minimum in $S$ for the warm Neptune scenario. This shows that it might be difficult to find a single bandpass combination (or observation strategy) that is convenient to differentiate Earth-like atmospheres from both warm and cold Neptunes.
In Figure 8 we show the significance of the various configurations of the two and three simultaneous bandpass scenarios (Options 2 and 3), for varying bandwidths. We notice a general increase in $S_{option}$ by considering two or three simultaneous bandpasses compared to the discovery bandpass only (D). This is by construction, since we selected the warm and cold Neptune models based on the smallest difference in flux in the discovery bandpass. The significance is very low in this case, but not zero, because of the uncertainty on the simulated measurement (condition 2, see Section IV).
@@
In the Warm Neptune case none of the options produce $S>3$, meaning that the difference in flux between the Warm Neptune and the Earth is less than three times the corresponding noise (as described in V). In the two-bandpass scenarios, Option 2a is generally preferred to the others for $\geq 10\%$ bandwidth ($S_{2a,10\%}=1.72$$S_{2a,20\%}=2.20$). When three bandpasses are available, Options 3a (UV+D+NIR) and 3c (UV+D+VIS) at 20% bandwidth are almost equally preferred, with $S\sim 2.7$.
@@
Disregarding for now the uncertainty associated with this measurement, we find that the planets are more effectively separated using narrow bandpasses: for the same option, the value of $d$ increases for decreasing bandwidths, resulting in points wider apart in color-color space. This has been observed also in Krissansen-Totton et al. (2016) for a similar study (see Section VII.3 for a more in depth comparison). However, the uncertainty on the flux measurements also becomes increasingly larger at smaller bandwidths (see Figure 6), which then factors into $\Delta d$. This often results in values of $r$ that are overall lower for smaller bandwidths. Exceptions arise when the bandpass combination changes with changes in bandwidth. We discuss this further in Section VII.2.
@@
From this analysis, we find that the combination of 360 nm + 500 nm + 1110 nm would be the most promising to differentiate both the cold and the warm Neptune case. However, the exposure time would need to be increased to reduce the uncertainties and obtain a value of $r\geq 3$; we discuss this in detail in Section VII.2.
VII Discussion
In this work, we assessed the potential of simultaneous multi-bandpass photometry as a key to breaking degeneracies inherent in single-bandpass, single-epoch observations of exoplanets. As shown in Sections III and IV, a single detection point can be explained by multiple orbital configurations, and planets with different compositions can produce similar fluxes in the single bandpass that HWO could use for its discovery phase (500 nm, 20% bandwidth). While longer integration times and improvements in the instrumentation can reduce uncertainties, they cannot fully eliminate this fundamental ambiguity.
Visiting the system at multiple epochs is the primary way to assess orbital parameters, and is therefore considered as a core part of HWO’s exoplanet survey strategy and taken into account when calculating exoplanet yields for the mission (see e.g., Stark et al., 2014; Morgan et al., 2024b). Such a strategy, however, comes with relevant complications: as time passes between each visit, the orbital position changes and therefore the portion of illuminated surface, the orientation of the planet, and the atmospheric dynamics would most likely have impacted the spectrum of the planet. Hence, multi-epoch observations are not trivial to interpret, and future work is still needed to model such variations, as well as accurately constrain the atmospheric composition. This is even more complicated for multi-planetary systems, where the pairing of the photometric detection over multiple epochs with the correct planet can be challenging and requires specialized algorithms (see e.g., Morgan et al., 2021; Hasler, 2025).
@@ VII Discussion
VII.1 Optimizing Bandpass Choices for Planet Differentiation
By using the Exposure Time Calculator that we developed (Section V, Alei et al. (2025a)) we estimate the measurement uncertainty for any potential ancillary bandpass given the observing time required to achieve S/NS/N=7 in the discovery observation (500 nm), the current baseline assumed by HWO. This allows us to identify the most promising bandpasses for various combinations of parallel channels to differentiate the Earth from “false positives” such as warm and cold Neptunes, thus informing an observational strategy for HWO optimized for multiple bandpasses (Section VI).
@@ VII.1 Optimizing Bandpass Choices for Planet Differentiation
This fundamental trade-off prevents the bandpass combinations from being universally optimal and highlights the model-dependent nature of our findings - we should expect that different models would yield different results and that a photometric observation strategy might take into account some additional population-related statistics (see Section VII.4).
@@
VII.2 $S/N$ Considerations
Figure 11: Total significance at varying $S/N$ values for the 2- (top row) and 3-bandpass options (bottom row) that were identified as promising to differentiate a warm or a cold Neptune from the Earth (see Section VI). In orange, solid lines the significance for the warm Neptune case; in aqua, dashed lines the significance for the cold Neptune case. As dashed lines, the $S_{option}=3$ and $S_{option}=5$ significance thresholds. Labeled in each subplots are the corresponding discovery exposure times (in hours) that are necessary to achieve $S_{option}=3$ and $S_{option}=5$, as calculated with pyEDITH. At the top of each subset, the option keyword, the corresponding bandwidth (BW) and the center wavelengths of the bandpasses considered in the option.
For the bandpass combinations that were identified as promising for discriminating the Earth from either Neptune-like atmosphere at the baseline discovery S/NS/N=7, we ran ancillary noise simulations assuming increasing discovery $S/N$ values. This is necessary to identify the bandpass combinations that would be best suited to qualitatively differentiate multiple planet types whenever possible, since we will not know a priori the nature of the planet when observing it for the first time. Hence, we analyzed how bandpasses that are optimized for a specific case would behave for the opposite scenario, for increasing observing time. This ancillary study provides information on the value that observing for longer exposure times has on the overall result over multiple scenarios, and allows us to prioritize specific bandpass combinations that favor both models at the same time, rather than a single scenario. We discuss future plans on how to generalize this effort in Section VII.4.
In Figure 11 we show the option significance that the two- and three-bandpass options we mentioned in the previous section yield for increasing discovery $S/N$ to differentiate both scenarios. We also label the exposure time that would be necessary to reach $S_{option}=3$ and $S_{option}=5$ significance levels for each model, calculated with pyEDITH. These times also include overheads to slew the telescope and dig the dark hole (see Section V).